OpenStudy (anonymous):

What value of r makes the equation true? r + 0.12 = 0.20 A. r = 0.32 B. r = 0.08 C. r = –0.08 D. r = –0.32

OpenStudy (anonymous):

You need to figure our what plus .12 = .20. The easiest way to do this is to do this. r + 0.12 = 0.20 Subtract 0.12. r = 0.08 now double check r + 0.12 = 0.20 0.08 + 0.12 = 0.20 0.20=0.20 Tada! :)

OpenStudy (anonymous):

Answer is B. :)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Have anymore? Your welcome. :)

OpenStudy (anonymous):

At 6:00 a.m., the outside temperature was –14⁰. By noon. the outside temperature had risen by 8⁰. What was the outside temperature at noon? A. –22⁰ B. –6⁰ C. 6⁰ D. 22⁰